Five divided by the difference of a number and 2 minus 4 divided by a number plus 2 equals 5 times the reciprocal of the difference of the number squared and 4. What is the​ number?
Answer by jorel1380(3719)

Let n be your number. Then:
5/(n-2) - 4/(n+2)=5/nē-4
5(n+2)-4(n-2)=5
5n+10-4n+8=5
n+18=5
n=-13 ☺☺☺☺